### v1

# Dynamic spillovers and investment strategies across artificial intelligence ETFs, artificial intelligence tokens, and green markets

Abstract
This paper investigates the risk spillovers among AI ETFs, AI tokens, and green markets using the $R^{2}$ decomposition method. We reveal several key insights. First, the overall transmission connectedness index (TCI) closely aligns with the contemporaneous TCI, while the lagged TCI is significantly lower. Second, AI ETFs and clean energy act as risk transmitters, whereas AI tokens and green bond function as risk receivers. Third, AI tokens are difficult to hedge and provide limited hedging ability compared to AI ETFs and green assets. However, multivariate portfolios effectively reduce AI tokens investment risk. Among them, the minimum correlation portfolio outperforms the minimum variance and minimum connectedness portfolios.
keywords:
AI ETFs , AI tokens , Green markets , Risk spillovers , Portfolio management
journal: Elsevier
1 Introduction
The onset of the Fourth Industrial Revolution has ushered human society into the era of artificial intelligence (AI). This transformation has not only revolutionized daily life, work practices, and communication methods but has also profoundly impacted education, healthcare, finance, and transportation [1]. Simultaneously, technological advancements have facilitated the evolution of financial markets, leading to the emergence of artificial intelligence related financial instruments. AI-related assets refer to a category of financial instruments closely associated with AI technology and its industrial ecosystem, encompassing AI-themed exchange-traded funds (AI ETFs), AI-related publicly traded stocks, and AI-integrated cryptocurrency assets (AI Tokens). AI ETFs allocate capital to companies in AI technology, robotics, cloud computing, and automation, offering diversified exposure to the AI sector [2]. AI-related stocks include companies where AI is a core business component or holds a significant role in revenue and strategic positioning, such as NVIDIA and Google [3, 4]. And AI Tokens integrate AI and blockchain technology, primarily facilitating decentralized AI computing, data sharing, and smart contract execution to support distributed AI development [3, 2]. These innovative asset classes not only represent the forefront of technological progress but have also become novel investment instruments, attracting significant interest from institutional and individual investors [2].
The interconnections among AI-related assets and and other instruments have also garnered significant attention, including stocks, bonds, commodities, cryptocurrencies, and electricity [2, 5, 4, 6, 7, 8, 9, 10, 11, 12]. Under normal market conditions, AI tokens can serve as effective diversifiers within portfolios containing traditional assets. However, during periods of extreme market stress, neither AI tokens nor AI ETFs provide significant risk diversification benefits for other asset classes. Moreover, AI tokens have the potential to provide cost-effective hedging for traditional assets [2]. The connectedness between AI assets and fossil fuel markets varies across time-frequencies and quantiles, with short-term fluctuations driving shock transmission, while long-term dynamics can shift assets between net transmitter and receiver roles [3]. The directional predictability between bitcoin, fintech, and AI stocks fluctuates across different time lags [6].
A growing body of research explores the interdependence among different types of AI-based assets [2, 9, 3, 13]. Directional effects indicate fluctuating connectedness across market indices, with AI stocks as persistent net transmitters, and AI tokens as primary shock recipients [9]. However, the correlation between AI stocks and tokens tends to be relatively low and may vary over time [13]. AI ETFs consistently act as net transmitters, dominating the system across market conditions, while AI tokens serve as weak net recipients at median quantiles, with all variables alternating between net transmitters and recipients over time [2]. SingularityNET (AGIX) and Fetch.ai (FET), two AI related tokens, consistently act as net recipients of volatility spillovers across all three frequency domains, while other markets assume varying roles depending on the investment horizon [3].
Some scholars have examined the determinants of spillovers and interactions between AI-related assets and other asset classes. Generally, the impact of the factors on spillover effects varies across different market conditions [4]. At the 5th and 95th percentiles, financial, economic, and uncertainty variables exhibit a highly similar impact on the spillover effects among AI-based assets, sukuk, and islamic equity indices [9]. Economic policy uncertainty and U.S. stock market volatility increase the interconnectedness between AI and fintech stocks, while geopolitical risk and gold market fluctuations have a negative impact on their correlations [11].
However, the rapid growth of AI technology, driven by deep learning and cloud computing, has sharply increased global energy demand. Its reliance on fossil fuels makes it a contributor to carbon emissions and a source of environmental concern [14]. Amid efforts to pursue a low-carbon economy, financial markets are increasingly integrating environmental factors into investment decisions and financing green projects, thus spurring the rise of green assets. Meanwhile, green assets also offer investors diversification and risk management opportunities [15]. Moreover, there exits potential connections between AI- related assets and sustainable markets. The empirical findings indicate that the AI index consistently emits shocks to clean markets across all frequencies, while the S&P Global Clean Energy Index functions as a net sender of volatility spillover across all market conditions and frequencies [7]. Investors incorporating stocks of artificial intelligence and robotics companies, cryptocurrencies, and green bonds into their portfolios should be aware that certain investment risks persist [10].
Although existing research has extensively explored the correlations between AI-related assets and other financial assets, studies on their interdependence with green markets remain relatively scarce. However, the development of green markets plays a crucial role in advancing the low-carbon economic transition and improving asset allocation efficiency. Furthermore, limited research has explored the role of sustainability-linked financial instruments in hedging risks and optimizing portfolio performance in relation to AI assets. Moreover, AI tokens, as a unique type of crypto asset, have introduced significant changes in global financial dynamics. However, research on their broader impact remains limited.
This study aims to quantify risk spillovers among AI ETFs, AI tokens, and green assets via the $R^{2}$ decompostion method [16, 17]. Further, we estimate the bilateral hedging ratios and hedge effectiveness of different assets pairs. Additionally, we apply the the minimum variance portfolio (MVP) [18], and compare its performance with the minimum correlation portfolio (MCP) [19], and the minimum connectedness portfolio (MCoP) [20]. This paper has the following contributions: First, examining risk spillovers among AI tokens, AI ETFs, and green assets helps uncover cross-market risk transmission patterns and potential contagion effects. This paper expands the existing literature by providing new insights into their interdependence. Second, by distinguishing between contemporaneous and lagged risk spillovers, this study provides a more comprehensive understanding of asset interdependencies over time. Third, the findings offer valuable insights for investors in optimizing asset allocation. Moreover, this study assists regulators in assessing market stability with a structured framework for risk transmission dynamics among AI and green assets.
The remainder of this paper is structured as follows. Section 2 provides a description of the data sets, along with summary statistics. Section 3 provides the methodological framework. Section 4 presents and discusses the empirical findings. Section 6 concludes the paper.
2 Data description and summary statistics
Throughout this research, we examine three categories of highly capitalized or actively traded assets: Firstly, AI ETFs, specifically iShares Robotics and Artificial Intelligence ETF (IRBO) and First Trust Nasdaq Artificial Intelligence and Robotics ETF (ROBT), which are representative assets in the artificial intelligence sector, with data sourced from Investing.com. Secondly, AI-based tokens, including Render (RENDER), NEAR Protocol (NEAR), Internet Computer (ICP), Filecoin (FIL), and Artificial Superintelligence Alliance (FET), obtained from CoinGecko. Among these tokens, FET was formed through the merger of Fetch.ai, SingularityNET, and Ocean Protocol, and was later rebranded as the Artificial Superintelligence token. In this study, it is referred to as FET for simplicity. These tokens represent diverse applications within the blockchain and AI ecosystems. RENDER focuses on decentralized graphics rendering. NEAR is dedicated to developing a scalable blockchain platform. ICP serves as the native token of the Internet Computer Protocol. FIL functions as the utility token for the Filecoin network, and FET is associated with an AI-powered decentralized machine learning network. Thirdly, green assets, comprising the S&P Green Bond (SPGB) and S&P Global Clean Energy Transition Index (SPGTCED), were collected from the S&P Global website. Due to data availability constraints, this study covers the period from May 25, 2021, to January 31, 2025, utilizing daily data with a total of 927 observations. All asset returns are calculated using logarithmic returns.
Table 1: Descriptive statistics.
Mean (×$\times 10^{2}$) Variance (×$\times 10^{2}$) Skewness Ex.Kurtosis JB ERS
IRBO 0.005 0.028 --0.003 1.717*** 113.810*** --12.922***
ROBT 0.002 0.024 0.069 1.264*** 62.427*** --14.157 ***
FET 0.494 0.677 1.252*** 7.151*** 2217.373*** --1.276 ***
FIL --0.054 0.474 1.738*** 17.811*** 12719.882*** --2.573***
ICP --0.054 0.504 1.352*** 10.751*** 4746.375*** --9.715 ***
NEAR 0.330 0.558 0.748*** 4.623*** 911.714*** --2.164 ***
RENDER 0.642 0.801 1.130*** 4.687*** 1045.970*** --1.498 ***
SPGB -0.019 0.023 0.247*** 1.825*** 138.021*** --6.615 ***
SPGTCED --0.061 0.025 0.454*** 2.307*** 237.196*** --14.460 ***
  • 1.
    Note: * , **, *** indicate significance at the 10%, 5%, and 1% levels, respectively.
Table 1 presents the descriptive statistics and preliminary tests for the AI and green assets. The mean returns vary significantly across assets, with RENDER and FET exhibiting the highest average returns, while FIL, ICP, SPGB, and SPGTCED display negative means. Variance suggests that AI tokens, particularly RENDER, exhibit higher volatility compared to AI ETFs and green assets. Skewness values indicate asymmetry in return distributions. While most assets exhibit positive skewness, IRBO shows slight letftward skew. Excess kurtosis values are significantly positive for all assets, implying leptokurtic distributions. The Jarque-Bera (JB) test strongly rejects normality for all assets at the 1% level. Additionally, the Elliott-Rothenberg-Stock (ERS) test results indicate strong stationarity for all series, as evidenced by significantly negative values. These findings suggest non-normal, highly volatile return distributions, particularly among AI tokens.
Figure 1: Pairwise correlation heatmap. Notes: Coefficients that do not meet the 10% significance level are shown as crossed-out values.
Fig. 1 presents a pairwise correlation heatmap for the selected assets, where the shading intensity reflects correlation strength. Darker red shades indicate higher positive correlations, while lighter shades correspond to weaker correlations. Correlation coefficients that do not meet the 10% significance level are marked with crossed-out values. The strongest correlation is observed between IRBO and ROBT (0.91), indicating a strong co-movement between these two AI-related ETFs. Among AI tokens, ICP and FIL (0.69) exhibit a relatively strong correlation, suggesting potential co-movement in their market dynamics, while NEAR and RENDER also display moderate correlations with other tokens, with values ranging from 0.43 to 0.58.
In contrast, the correlations between SPGB (green bonds) and AI assets remain relatively weak, with values mostly below 0.3, suggesting limited interaction between these asset classes. However, SPGTCED (another green asset) shows a stronger correlation with AI assets, reaching 0.62 with ROBT and 0.58 with IRBO, suggesting some level of connectedness between sustainable investment themes and AI-driven assets. Notably, The correlation coefficients between ROBT and AI tokens, as well as those between clean energy assets and cryptos, largely fail to pass the significance test.
In conclusion, the results indicate strong intra-category correlations. Green assets exhibit moderate correlations with AI ETFs. However, AI tokens have weak or insignificant correlations with other asset classes.
3 Methodology
3.1 $R^{2}$ decomposition approach
In this study, we utilize the innovative $R^{2}$ decomposition technique to examine the dynamic linkages among AI-based ETFs, Tokens, and green markets over different time periods [16, 17]. This approach efficiently disaggregates relevance into contemporaneous and lagged components while ensuring standardized data processing. By overcoming the limitations of prior research, which predominantly concentrated on contemporaneous spillover effects, this method broadens the scope of investigation [16]. The total connectedness index (TCI) is computed as the average $R^{2}$ across $k$ multivariate linear regressions:
$TCI=\frac{1}{K}\sum_{k=1}^{K}R_{k}^{2}=\left(\frac{1}{K}\sum_{k=1}^{K}\sum_{i=1}^{K}\bm{R}_{C,k,i}^{2,d}\right)+\left(\frac{1}{K}\sum_{k=1}^{K}\sum_{i=1}^{K}\bm{R}_{L,k,i}^{2,d}\right)=TCI^{C}+TCI^{L},$ (1)
where $\bm{R}^{2,d}_{C}$ and $\bm{R}^{2,d}_{L}$ refer to contemporaneous and lagged spillover effects, respectively. The total directional spillovers, both to and from other markets, as well as the net spillover, are calculated as follows:
$TO_{i}$ $=\sum_{k=1}^{K}\bm{R}_{C,k,i}^{2,d}+\sum_{k=1}^{K}\bm{R}_{L,k,i}^{2,d}=TO_{i}^{C}+TO_{i}^{L},$ (2)
$FROM_{i}$ $=\sum_{k=1}^{K}\bm{R}_{C,i,k}^{2,d}+\sum_{k=1}^{K}\bm{R}_{L,i,k}^{2,d}=FROM_{i}^{C}+FROM_{i}^{L},$ (3)
$NET_{i}$ $=TO_{i}-FROM_{i}$ (4)
$=(TO_{i}^{C}+TO_{i}^{L})-(FROM_{i}^{C}+FROM_{i}^{L})$ (5)
$=(TO_{i}^{C}-FROM_{i}^{C})+(TO_{i}^{L}-FROM_{i}^{L})$ (6)
$=NET^{C}_{i}+NET^{L}_{i}.$ (7)
When $NET_{i}>0$ (or $NET_{i}<0$), the market $i$ is identified as a net contributor (or receiver) of shocks. Furthermore, net pairwise directional connectedness (NPDC) is calculated in a similar manner to NET, but on a pairwise level [16].
3.2 Portfolio analysis
3.2.1 Bilateral hedge ratios and portfolio weights
According to Kroner et al. [21], the dynamic hedge ratio is expressed as:
$\beta_{ij,t}=\frac{\Sigma_{ij,t}}{\Sigma_{jj,t}},$ (8)
where $\Sigma_{ij,t}$ and $\Sigma_{jj,t}$ represents the conditional covariance and the conditional variance at time $t$.
Following Kroner and Ng [22], the optimal bilateral portfolio weights are given by:
$w_{ij,t}=\frac{\Sigma_{ii,t}-\Sigma_{ij,t}}{\Sigma_{ii,t}-2\Sigma_{ij,t}+\Sigma_{jj,t}},$ (9)
with
$w_{ij,t}=\begin{cases}0,&\text{if }w_{ij,t}<0\\ w_{ij,t},&\text{if }0\leq w_{ij,t}\leq 1,\\ 1,&\text{if }w_{ij,t}>1\end{cases}$ (10)
where $w_{ij,t}$ denotes the proportion of series $i$ in a $1$-dollar portfolio consisting of series $i$ and $j$ at time $t$. Consequently, $1-w_{ij,t}$ represents the weight of series $j$ within the same portfolio.
3.2.2 Minimum variance portfolio
The minimum variance portfolio (MVP) forms lowest possible risk for a set of assets [18], which is widely used in portfolio management to construct a low-volatility portfolio. The optimal weights for MVP are given by:
$\mathbf{w}_{\Sigma_{t}}=\frac{\Sigma_{t}^{-1}I}{I^{T}\Sigma_{t}^{-1}I},$ (11)
where $\mathbf{w}_{\Sigma_{t}}$ represents the portfolio weight vector, $I$ is a $K$-dimensional vector of ones,and $\Sigma_{t}$ denotes the $K\times K$ conditional variance-covariance matrix at time $t$.
3.2.3 Minimum correlation portfolio
The minimum correlation portfolio (MCP) method is another widely used portfolio construction approach, introduced by Christoffersen et al. [19]. This method constructs portfolios by minimizing conditional correlation. The portfolio weights under MCP are computed as follows:
$R_{t}=\text{diag}(\Sigma_{t})^{-0.5}H_{t}\text{diag}(\Sigma_{t})^{-0.5},$ (12)
where $R_{t}$ is the conditional correlation matrix for the $K\times K$ dimensional portfolio at time $t$. The weights in MCP are calculated as:
$\omega_{R_{t}}=\frac{R_{t}^{-1}I}{I^{T}R_{t}^{-1}I},$ (13)
where $\omega_{R_{t}}$ is the weight vector for the $K\times 1$ dimensional portfolio, and $I$ represents the $K$-dimensional vector of ones.
3.2.4 Minimum Connectedness Portfolio
The minimum connectedness portfolio (MCoP) approach constructs a portfolio by minimizing the pairwise connectedness of indices [20]. The asset weights are computed as follows:
$\omega_{C_{t}}=\frac{PCI_{t}^{-1}I}{IPCI_{t}^{-1}I},$ (14)
where $\omega_{C_{t}}$ represents the weight of an asset in the portfolio, $PCI_{t}$ denotes the pairwise connectedness index matrix at time $t$, and $I$ is the identity matrix.
3.2.5 Portfolio analysis
To evaluate portfolio performance, two key metrics are used: the Sharpe ratio (SR) [23] and hedging effectiveness (HE) [24].
The Sharpe ratio is computed as:
$SR=\frac{\bar{r}_{p}}{\sqrt{\text{var}(r_{p})}},$ (15)
where $r_{p}$ represents portfolio returns assuming a risk-free rate of zero. A higher SR indicates superior returns relative to portfolio risk, facilitating comparison across portfolios with similar volatility.
The hedging effectiveness (HE) measures the percentage risk reduction achieved by the portfolio compared to investing in a single asset $i$ [24]. It is given by:
$HE_{i}=1-\frac{\text{var}(r_{p})}{\text{var}(r_{i})},$ (16)
where $\text{var}(r_{p})$ denotes the portfolio variance and $\text{var}(r_{i})$ represents the variance of asset $i$. A higher HE value signifies greater risk reduction.
4 Dynamic connectedness analysis
4.1 Averaged spillover effects
As illustrated in Fig. 2, the left panel of the figure presents the heatmaps of the average overall, contemporaneous, and lagged connectness. The $(i,j)$ element represents the average spillover from asset $j$ to asset $i$, while the $(j,i)$ element denotes the average spillover from asset $i$ to asset $j$. The difference between these values reflects the net average spillover from asset $j$ to asset $i$. The three heatmaps suggest that the level of information transmission among different types of AI assets remains relatively moderate in both TO and FROM connectedness, mirroring the weak correlation between AI stocks and AI tokens [13]. However, the spillover effects within the same asset are relatively stronger.
Furthermore, AI ETFs act as major net spillover transmitters to both green bonds and AI tokens, a finding that aligns with the results reported by Yousaf et al. [2] and Huynh et al. [10]. A more detailed examination reveals that ROBT transmits net spillovers to all assets, including IRBO and clean energy. This result is consistent with Zeng et al.’s work [7], which highlights that the NASDAQ CTA Artificial Intelligence & Robotics Index, tracked by ROBT, also transmits risk to clean energy. The spillover dynamics of AI tokens exhibit strong similarities, potentially driven by the herding effect in the artificial intelligence and big data token markets. Investors appear to be influenced by collective behavior, leading to coordinated market movements that align with prevailing market consensus [25].
Figure 2: Averaged dynamic spillover indices. Notes: In the heatmap, the $(i,j)$ element represents the magnitude of the spillover effect from asset $j$ to asset $i$. Panels (a) through (f) correspond to overall, contemporaneous, lagged, TO, FROM, and NET connectedness, respectively.
A closer examination indicates that the overall and contemporaneous average spillover effects exhibit similar patterns. First, within the same asset category, spillover effects between assets tend to be stronger. For instance, the overall spillovers from IRBO to ROBT and from ROBT to IRBO are 62.37 and 61.14, respectively, indicating the strongest risk transmission between these two assets. This result is consistent with their highest positive correlation, as shown in Fig. 1. A similar pattern is observed in average contemporaneous connectness. Second, within the AI token market, FIL and ICP exhibit the strongest mutual spillover transmission. Third, spillover effects between clean energy assets and AI ETFs are also relatively strong, indicating notable risk transmission between these asset classes.
However, a closer examination of Fig. 2 (a-c) reveals distinctions among the averaged overall, contemporaneous, and lagged spillover effects. Specifically, in the overall analysis, clean energy transmits net spillovers to all assets, including IRBO and bonds, except for ROBT. In the contemporaneous period, all assets transmit net spillovers to green bond. In the lagged period, clean energy transmits spillovers to tokens and bonds.
Observing Fig. 2 (d-f), the three bar charts on the right illustrate the averaged spillover effects among different assets, with each panel representing overall (black), contemporaneous (red), and lagged (green) connectedness. Fig. 2 (d) shows that AI ETFs (IRBO and ROBT) and clean energy have the highest TO connectenss, indicating that they are the primary risk transmitters, with most spillover effects occurring contemporaneously rather than with a lag. Fig. 2 (e) presents FROM connectedness, measuring the extent to which each asset receives risk from others. AI ETFs and clean energy continue to exhibit the most significant FROM values, suggesting that they not only transmit risk but also absorb it from other assets. Lagged effects remain relatively small, indicating that risk transmission is predominantly immediate. Fig. 2 (f) represents NET connectedness, calculated as the difference between TO and FROM spillovers. AI ETFs have the highest net spillover values, reinforcing their role as key risk transmitters, while AI tokens and green bond generally exhibit negative net spillover values, confirming their role as risk recipients. Notably, SPGTCED has a positive net spillover value, suggesting that it still transmits a certain level of risk to other assets. All assets display lagged spillover effects, indicating that NET risk transmission may experience slight delays over time. Overall, these charts suggest that AI ETFs and clean energy serve as the primary sources of risk transmission, whereas AI tokens and green bonds mainly act as risk recipients, with NET spillover effects occurring primarily contemporaneously rather than over extended periods. The empirical results are in line with work of Billah [9], Yousaf et al. [3], and Yousaf et al. [2].
4.2 Dynamic total connectedness measures
Fig. 3 presents the dynamic total connectedness, where both overall and contemporaneous connectedness exhibit a similar upward trend. The first peak occurs in mid-December 2022, reaching approximately 64%, followed by a decline. Throughout 2023, some fluctuations are observed. The second major peak occurs in mid-January 2025, reaching approximately 66%. In contrast, lagged TCI remains consistently low, fluctuating below 10% throughout the entire period.
Figure 3: Evolution of total connectedness.
Overall, the strong synchronization between overall and contemporaneous connectedness suggests that risk spillovers primarily occur in real-time rather than with a lag. The consistently low lagged connectedness indicates that market information transmission does not experience significant delays.
4.3 NET total directional connectedness
Fig. 4 illustrates the time-varying directional connectedness of AI and green assets. From the figure, it can be observed that the overall spillover effects and lagged spillover effects exhibit a closely aligned trend across various asset classes. In contrast, the net contemporaneous spillover effects show relatively weaker fluctuations, with certain assets experiencing significant increases in spillover effects during specific periods.
Figure 4: NET total directional connectedness.
The AI ETFs and clean energy assets exhibit a similar trend, both reaching a peak around October 2022, followed by a decline, and then an upward trajectory starting in 2024. Their overall spillovers remain predominantly positive throughout the period, consistently serving as the strongest contributors to systemic risk, highlighting their role as key drivers of market fluctuations. Ref. [7] and this study both conclude that clean energy acts as a net transmitter. Additionally, while Zeng et al. identify the AI index as a net transmitter [7], this study finds that ROBT tracking the AI index also serves as a net transmitter. Consistent with Billah’s findings [9], AI tokens likewise act as recipients of shocks within the system.
In contrast, AI tokens and SPGB maintain relatively low levels of overall spillover dynamics, which remain negative for most of the period, except between mid-2023 and mid-2024. Their behavior is almost the opposite of AI ETFs and clean energy assets, indicating that they generally act as recipients of shocks within the system. This is particularly evident during periods of heightened market volatility, where their net spillover inflows become more pronounced. However, it is important to note that the spillover roles of these assets evolve dynamically over time, and in certain periods, they may also serve as risk transmitters.
Overall, Fig. 4 reflects the complex interrelationships between the AI and green asset markets. AI ETFs and clean energy assets exhibit strong similarities, whereas AI tokens and the green bond market follow distinct spillover effect patterns. This suggests that different types of AI and green assets interact differently with systemic risk, highlighting the necessity of further exploration into their interconnectedness under varying market conditions.
4.4 Net pairwise directional connectedness measures
Fig. 5 illustrates the net pairwise directional spillover. The dynamics of risk transmission from these two ETFs to other assets are highly similar, and ROBT primarily transmits risk to IRBO. Aligning with the findings of Huynh et al. [10], AI funds exhibits high overall spillovers, which is consistent with Ref.  [2]. Moreover, the spillover effects between AI ETFs and AI tokens exhibit a similar pattern, generally increasing throughout the sample period until reaching a peak, followed by a decline, reaching their lowest levels between late 2023 and the first quarter of 2024, before rising again.
Figure 5: Net pairwise directional connectedness.
Besides that, there exists significant spillovers from AI based ETFs to green bond, in line with Fig. 2. The connectedness declines until the end of 2023 and then stabilizes, fluctuating around zero. The spillover from AI ETFs to green bonds is generally stronger than the spillover from AI ETFs to clean energy. In addition, all AI tokens receive spillovers from clean energy with similar dynamics. It first reaches its maximum value in the second half of 2022, then declines and fluctuates around zero. NEAR and clean energy are exceptions, as their connectedness reaches its strongest level in the first quarter of 2023. Overall, assets of the same type generally alternate between the roles of transmitter and receiver. AI ETFs and clean energy mainly acts as contributors of system shocks to other assets, which accords with Fig. 4.
4.5 Net pairwise network
Fig. 6 depicts the net pairwise directional spillover networks, filtering edges below 0.05 to highlight major spillover effects. Blue nodes represent transmitters, while yellow nodes denote receivers. In line with the results documented in Ref. [15], our study also identifies a risk spillover effect from clean energy to green bonds. Again, the AI ETFs is the primary shock transmitters to other assets. Connectedness among AI tokens are weaker than their links with AI ETFs and green assets. The overall and lagged networks suggests the strong spillovers from AI ETFs and clean energy to other assets. ROBT is the largest transmitter, followed by IRBO, and finally clean energy. Yousaf et al. also report that AI ETFs (tokens) are strong (weakest) net transmitters (recipients) during normal market conditions [2]. AI tokens and green bond receive relatively similar level of shocks from other markets. However, the contemporaneous network indicates that IRBO, FIL, RENDER, and SPGB acts as receivers. Green bond receives most shocks from other markets, followed by IRBO. Further, FET, ICP, and NEAR transmit risk, although the level of shock remains relatively low.
Figure 6: Net pairwise directional spillover networks.
4.6 Robustness test
As shown in Fig. 7, we conduct a robustness analysis to verify the reliability of our findings, employing multiple approaches, including Pearson, Spearman, and Kendall correlation-based $R^{2}$ decompositions, as well as quantile VAR (QVAR) [26]. All TCIs exhibit a largely consistent pattern, reinforcing the robustness of our findings.
Figure 7: Robustness analysis for TCI.
5 Portfolio and hedging strategies
5.1 Bilateral hedge ratios and portfolio weights
Table 2 presents the summary statistics of bilateral hedge ratios and hedge effectiveness (HE) for AI and green assets. First, AI ETFs and green assets demonstrate relatively strong hedging effectiveness in both long and short positions. Their mutual hedging performance is also robust. ROBT/SPGB (0.957) and IRBO/SPGB (0.946) show the highest hedge ratios, both statistically significant ($p$-value $<$ 0.01). When hedging AI ETFs, clean energy outperforms green bonds in both long and short positions.
Table 2: Bilateral hedge ratios and hedge effectiveness.
      Long/Short       Mean       Std. Dev.       5%       95%       HE       p-value
      IRBO/FET       0.007       0.011       --0.010       0.022       0.018       0.886
      IRBO/FIL       0.008       0.017       --0.018       0.041       0.018       0.855
      IRBO/ICP       0.000       0.017       --0.024       0.037       0.014       0.887
      IRBO/NEAR       0.009       0.014       --0.020       0.030       0.018       0.850
      IRBO/RENDER       0.012       0.010       --0.002       0.029       0.020       0.840
      IRBO/SPGB       0.946       0.334       0.284       1.486       0.102       0.002
      IRBO/SPGTCED       0.626       0.121       0.304       0.778       0.372       0.000
      ROBT/FET       0.004       0.011       --0.013       0.022       0.018       0.886
      ROBT/FIL       0.000       0.017       --0.028       0.028       0.017       0.855
      ROBT/ICP       --0.008       0.017       --0.033       0.027       0.012       0.887
      ROBT/NEAR       0.004       0.013       --0.017       0.024       0.018       0.850
      ROBT/RENDER       0.004       0.011       --0.013       0.023       0.017       0.840
      ROBT/SPGB       0.957       0.327       0.388       1.516       0.126       0.002
      ROBT/SPGTCED       0.609       0.114       0.367       0.776       0.412       0.000
      FET/IRBO       0.185       0.285       --0.225       0.772       0.016       0.000
      FET/ROBT       0.106       0.291       --0.316       0.571       0.015       0.000
      FET/SPGB       --0.520       1.425       --3.129       2.175       0.008       0.002
      FET/SPGTCED       --0.104       0.284       --0.570       0.326       0.009       0.000
      FIL/IRBO       0.139       0.277       --0.261       0.632       0.015       0.000
      FIL/ROBT       --0.015       0.283       --0.468       0.390       0.015       0.000
      FIL/SPGB       0.418       0.830       --0.677       1.963       0.016       0.002
      FIL/SPGTCED       --0.027       0.274       --0.466       0.486       0.012       0.000
      ICP/IRBO       0.035       0.302       --0.421       0.713       0.013       0.000
      ICP/ROBT       --0.120       0.306       --0.527       0.543       0.011       0.000
      ICP/SPGB       --0.165       0.715       --1.623       1.093       0.003       0.002
      ICP/SPGTCED       --0.091       0.266       --0.422       0.462       0.009       0.000
      NEAR/IRBO       0.115       0.375       --0.880       0.624       0.018       0.000
      NEAR/ROBT       0.066       0.327       --0.664       0.503       0.017       0.000
      NEAR/SPGB       0.697       0.906       --1.004       2.157       0.015       0.002
      NEAR/SPGTCED       --0.121       0.425       --1.306       0.373       0.012       0.000
      RENDER/IRBO       0.295       0.250       --0.126       0.699       0.018       0.000
      RENDER/ROBT       0.108       0.331       --0.400       0.777       0.015       0.000
      RENDER/SPGB       0.015       0.971       --1.415       1.618       0.009       0.002
      RENDER/SPGTCED       --0.079       0.305       --0.531       0.434       0.013       0.000
      SPGB/IRBO       0.088       0.044       0.018       0.150       0.118       0.000
      SPGB/ROBT       0.098       0.044       0.027       0.163       0.134       0.000
      SPGB/FET       --0.002       0.004       --0.009       0.006       0.010       0.886
      SPGB/FIL       0.002       0.004       --0.003       0.009       0.015       0.855
      SPGB/ICP       0.000       0.003       --0.005       0.005       0.007       0.887
      SPGB/NEAR       0.004       0.004       --0.002       0.011       0.016       0.850
      SPGB/RENDER       0.001       0.003       --0.003       0.006       0.011       0.840
      SPGTCED/IRBO       0.605       0.118       0.266       0.749       0.383       0.000
      SPGTCED/ROBT       0.661       0.116       0.330       0.807       0.414       0.000
      SPGTCED/FET       --0.004       0.010       --0.022       0.010       0.011       0.886
      SPGTCED/FIL       0.000       0.015       --0.025       0.028       0.014       0.855
      SPGTCED/ICP       --0.009       0.016       --0.036       0.020       0.008       0.887
      SPGTCED/NEAR       --0.003       0.014       --0.034       0.018       0.011       0.850
      SPGTCED/RENDER       --0.002       0.011       --0.020       0.016       0.013       0.840
Second, AI tokens provide limited effectiveness when used to hedge AI ETFs. Hedge ratio means are near zero, HE remains low (0.011-0.020), and $p$-values exceed 0.8, indicating poor risk reduction capabilities. When tokens are long and AI ETFs are short, hedge ratios increase. The highest hedge ratio is observed in RENDER/IRBO (0.295), while NEAR/IRBO achieves the highest HE (0.018), still providing minimal risk mitigation. Compared to Ref. [2], the hedge ratio between AI ETFs and tokens is highly similar. However, the hedging effectiveness differ. This discrepancy may stem from differences in the time period analyzed and sample selection.
Third, green assets fail to provide a strong safe-haven function against cryptos. FIL/SPGB (HE = 0.016, $p$-value $<$ 0.01) shows the best performance, but still offers weak risk reduction. When green assets are in a long position, the results lack statistical significance. The HE for all tokens remains no bigger than 0.016 (SPGB/NEAR), indicating the weak hedging ability of AI tokens.
5.2 Multivariate portfolio analysis
Table 3 show the results for MVP, MCP and MCoP. In the MVP strategy, the relatively high HE values (IRBO: 0.905, ROBT: 0.890) indicate that multivariate portfolios are effective in reducing the risk of AI ETFs. Even though AI ETFs are the primary transmitters of shocks in the return transmission analysis, their weights remain relatively low (IRBO: 2.4%, ROBT: 3.8%). This suggests that their role in the MVP portfolio is limited. In the MCP, the allocation to AI ETFs increases significantly (IRBO: 6.0%, ROBT: 12.5%), indicating an enhanced contribution to the portfolio. However, their HE values turn negative (IRBO: $-2.215$, ROBT: $-2.728$), indicating that this portfolio fails to effectively mitigate volatility in AI ETFs and may even increase risk. In the MCoP method, the allocation to AI ETFs further increases (IRBO: 10.5%, ROBT: 11.5%), reinforcing their role within the portfolio. Nevertheless, their HE values continue to decline (IRBO: $-3.375$, ROBT: $-4.073$). Overall, despite Figs. 2, 4 and 6 indicating that AI ETFs are major risk spillovers, their weights in these portfolios remain relatively limited.
Next, we analyze the role of AI Tokens across different portfolio strategies. In the MVP, AI tokens (FET, FIL, ICP, NEAR, RENDER) have minimal weights, each below 1%, limiting their overall impact on the portfolio. Furthermore, their HE values exceed 0.99, suggesting that this investment strategy effectively minimizes their risk. In the MCP portfolio, despite a decrease in HE values (ranging between 0.724 and 0.854), the weights of AI Tokens increase significantly (e.g., RENDER: 12.6%, FET: 11.6%), indicating their enhanced role within the portfolio. In the MCoP, the weights of AI Tokens further increase, reflecting their rising importance in this investment strategy. Their HE values remain above 0.7 (ranging between 0.705 and 0.802), suggesting that this strategy contribute to volatility reduction of AI tokens, though to a lesser extent than in the MVP. Overall, AI Tokens exhibit higher allocation ratios for both MCP and MCoP, making them core assets, whereas their allocation in the MVP remains low due to higher volatility, limiting their overall influence.
Finally, we examine the performance of green assets in the portfolios. Green assets hold the highest allocation in the MVP, particularly green bonds (SPGB: 91.2%, SPGTCED: 0.7%). Green assets remain the core holdings in the MVP, with the highest allocation. SPGB holds an overwhelming weight of 91.2%, while SPGTCED accounts for 0.7%, likely reflecting their respective roles as a risk receiver and a risk contributor. Accordingly, SPGTCED has an HE value of 0.884, demonstrating a significantly stronger risk reduction effect compared to SPGB, which has a negative HE value of --0.05. From the MCP procedure, the weight of green assets declines significantly (SPGB: 19.7%, SPGTCED: 13.2%, total: 32.9%), and their HE values continue to decline, indicating a weakening risk reduction in this portfolio. In the MCoP, the allocation to green assets decreases further (SPGB: 11.2%, SPGTCED: 11.0%, total: 22.2%), with HE values continuing to drop, suggesting a further reduction in their role. Despite Figures 2, 4 and 6 demonstrating that clean energy transmits shocks to other markets, its weight in the portfolio is not dominant. This finding aligns with previous literature [15], which also found that while clean energy is a risk contributor, its allocation in investment portfolios is not necessarily the highest, whereas green bonds tend to account for a larger proportion of portfolio allocation.
Table 3: Multivariate Portfolio Weights.
Mean Std.Dev. 5% 95% HE p-value
MVP IRBO 0.024 0.028 0.000 0.072 0.905 0.000
ROBT 0.038 0.044 0.000 0.121 0.890 0.000
FET 0.007 0.006 0.000 0.016 0.996 0.000
FIL 0.003 0.003 0.000 0.009 0.994 0.000
ICP 0.006 0.005 0.000 0.015 0.994 0.000
NEAR 0.000 0.001 0.000 0.000 0.994 0.000
RENDER 0.002 0.002 0.000 0.007 0.996 0.000
SPGB 0.912 0.037 0.846 0.970 --0.050 0.508
SPGTCED 0.007 0.014 0.000 0.043 0.884 0.000
MCP IRBO 0.060 0.082 0.000 0.240 --2.215 0.000
ROBT 0.125 0.087 0.000 0.232 --2.728 0.000
FET 0.116 0.047 0.028 0.183 0.854 0.000
FIL 0.080 0.051 0.007 0.170 0.792 0.000
ICP 0.113 0.064 0.000 0.207 0.783 0.000
NEAR 0.052 0.042 0.000 0.124 0.785 0.000
RENDER 0.126 0.027 0.086 0.170 0.851 0.000
SPGB 0.197 0.034 0.139 0.261 --34.611 0.000
SPGTCED 0.132 0.034 0.061 0.176 --2.927 0.000
MCoP IRBO 0.105 0.112 0.000 0.250 --3.375 0.000
ROBT 0.115 0.120 0.000 0.250 --4.073 0.000
FET 0.104 0.115 0.000 0.250 0.802 0.000
FIL 0.109 0.111 0.000 0.250 0.718 0.000
ICP 0.114 0.119 0.000 0.250 0.705 0.000
NEAR 0.111 0.116 0.000 0.250 0.708 0.000
RENDER 0.120 0.128 0.000 0.333 0.798 0.000
SPGB 0.112 0.118 0.000 0.250 --47.457 0.000
SPGTCED 0.110 0.113 0.000 0.250 --4.343 0.000
Interestingly, in the case of the MVP, asset allocations, except for AI tokens, are directly opposite to their roles. Risk receivers have relatively higher weights, while risk transmitters have lower weights. The results of MCP and MCoP are quite similar, possibly because both focus on reducing correlations across assets. Both portfolios reduce the risk of AI tokens but increase the volatility of the other two asset classes.
5.3 Portfolio performance
We illustrate the cumulative portfolio returns in Fig. 8. This figure presents the cumulative returns of three portfolio strategies: MVP (black line), MCP (red line), and MCoP (light blue line). MVP remains relatively stable with minimal fluctuations, indicating a low-risk but low-return profile, while MCP and MCoP show significant growth, especially from early 2023 onward. In the early period, all portfolios experience negative returns, but MCP and MCoP recover strongly and reach cumulative returns above 1.0 by early 2025, whereas MVP remains around the zero level. MCP and MCoP follow similar trends, with MCoP generally achieving slightly higher peaks, suggesting better performance in specific market conditions. Other studies on investment strategy comparisons also documents negative cumulative portfolio returns [27, 28].
Figure 8: Cumulative returns of different portfolios.
Finally, the Sharpe ratios of the three portfolios are shown in Table 4. The MCP emerges as the most optimal strategy, as it achieves the highest return and Sharpe ratios across all evaluated metrics. While the MVP effectively minimizes volatility, its negative return and Sharpe ratios render it an unattractive investment option. The MCoP offers a moderate risk-return profile, yet it fails to surpass MCP. Therefore, for investors seeking an optimal balance between risk and return, the MCP portfolio represents the most advantageous selection. This finding is similar with Ziadat et al. [29], who also observe a negative Sharpe ratio for the MVP, while the MCP outperforms both the MVP and MCoP.
Table 4: Portfolio performance.
           MVP            MCP            MCoP
           Return            --0.0165            0.2664            0.2438
           StdDev            0.0835            0.4860            0.5669
           Sharpe Ratio (StdDev)            --0.1978            0.5482            0.4300
           Sharpe Ratio (VaR)            --1.9865            5.8477            4.8230
           Sharpe Ratio (CVaR)            --1.4889            4.2027            3.8360
6 Conclusion
This study employs the $R^{2}$ decomposition to examine the risk spillover dynamics among AI ETFs, AI tokens, and green assets over the period from May 25, 2021, to January 31, 2025. Our findings reveal that contemporaneous spillovers dominate the total connectedness index, while lagged spillovers remain relatively stable. Furthermore, AI ETFs and clean energy emerge as risk transmitters, whereas AI tokens and green bonds primarily act as risk receivers. In addition, AI tokens are challenging to hedge and offer little benefit in hedging other assets when considering bilateral hedge ratios. However, clean energy and AI ETFs exhibit stronger hedging effectiveness regardless of the long or short position taken. Additionally, we analyze the performance of three portfolio strategies: MVP, MCP, and MCoP. The results show that all the approaches reduce volatility of AI tokens significantly. However, only the MVP lowers the risk of other assets, while the other two strategies actually increase their risk. Finally, MCP exhibits superior performance than the other portfolios. In this investment strategy, green bonds have the largest allocation, while IRBO has the smallest.
Our work has the following implications: First, Investors should consider the high spillover risks associated with AI ETFs and clean energy. Although the two assets classes are highly popular, their risk transmission role suggests that they should be complemented with strong hedging assets. Besides that, given the ineffective hedging of AI tokens with other assets, investors should closely monitor their exposure to AI-related cryptos and adjust their allocations accordingly. MCP outperforms MVP and MCoP, making it the most suitable strategy for investors, especially those interested in AI tokens. Second, given the risk-spreading nature of AI ETFs and clean energy assets, regulators should closely monitor their volatility and systemic impact. Implementing safeguards to prevent excessive market fluctuations is essential. Additionally, policymakers should support financial instruments that facilitate the integration of these assets into broader portfolios, preventing them from becoming hidden sources of risk.